OpenStudy (anonymous):

Solve –4x + 7 = x + 2

OpenStudy (anonymous):

-4x + 7 = x + 2 Add 4x and subtract 2 5 = 5x Divide by 5 1 = x Done.

OpenStudy (anonymous):

thnx i see now